# Disorder to Order Transition in 1D Nonreciprocal Cahn-Hilliard Model

### v1

@@ III.2 Wave number selection
A we show in Fig. 3, for a given $\alpha$, sources and sinks with a particular value of the asymptotic wave number $k_{\infty}$ are selected. $k_{\infty}$ increases monotonically with $\alpha$, and from a least-square fit, we obtain $k_{\infty}\sim\alpha^{0.6}$. Further, $k_{\infty}$ is comparable to the wave number selected for targets in 2D. In Fig. 3(inset), we verify that $k_{\infty}$ for a single source-sink pair configuration is the same as the $k_{\infty}$ obtained from multi-defect configurations that evolved from disordered initial conditions. A monotonically increasing $k_{\infty}$ predicts a crossover value of non reciprocity, $\alpha_{\times}$, such that when $\alpha\to\alpha_{\times}$, $k_{\infty}\to 1$, and $R_{\infty}\to 0$. It implies that the defect solutions will cease to exist for $\alpha\geq\alpha_{\times}$. Using a least-square fit, we find $\alpha_{\times}\sim 1.5$. However this prediction of $\alpha_{\times}$ is not entirely correct as it does not account for the Eckhaus instability. Since the plane waves generated by the defects are unstable for $k^{2}\geq 1/3$ [23, 8, 3, 29], we expect that the defect solutions will vanish for $k_{\infty}^{2}\geq 1/3$ which yields $\alpha_{\times}\sim 0.62$. Thus wavenumber selection and the Eckhaus instability restrict the defect solutions to the range $0<\alpha<0.62$.
@@ III.3 Defect density and the Global polar order
We now focus on the statistical properties of the multi-defect configurations. Using the fact that $J=0$ at a source or a sink, we define the 1D defect density $\rho_{D}$ simply as the number of zeros of $J$ per unit length. In Fig. 4(a), we plot $\rho_{D}$ in the steady state for different values of $\alpha$. Since $\alpha=0$ is the equilibrium limit of the model, $\rho_{D}$ vanishes as $\alpha\to 0$. Furthermore, we find that defects cease to exist above $\alpha_{c}\sim 0.6$ and consequently $\rho_{D}$ vanishes, instead we find travelling waves. This marks the onset of disorder-to-order transition that was also observed in 2D [3]. As shown in Fig. 4(b), we plot the average polar order $\bar{J}$ for different values of $\alpha$, in the steady state. $\bar{J}$ is defined as
@@
where $\left$$\ldots\right$> implies averaging over space and time in the steady-state, and we use the dominant wave number of the patterns $q_{p}$ for normalization. For travelling waves, $J=R^{2}k=k-k^{3}$, and $q_{p}=k$, which sets $\bar{J}=1$ for all travelling waves, irrespective of their wave number. For defects, $q_{p}=k_{\infty}$ is a suitable choice. Since defects are sources or sinks of travelling waves in both left and right directions, we expect $\bar{J}$ to vanish in the presence of defects. For small $\alpha$, $\bar{J}$ is close to zero. At large $\alpha$, consistent with travelling states, $\bar{J}\sim 1$. In the vicinity of the transition point $\alpha_{c}$ (grey shaded region), $\bar{J}$ shows large fluctuations as some of the simulations can transition to travelling wave states while others show defects. We verify this by plotting the fraction of simulations $(F_{T})$ that show defect states for $\alpha$ in the shaded region (see Fig. 4(b,inset)). For $\alpha\lesssim 0.4$, the entire ensemble shows defect states whereas for $\alpha>0.6$, all the simulations end up in travelling wave states. From the plots of $\rho_{D}$ and $\bar{J}$, we conclude that the disorder to order transition in 1D occurs at $\alpha_{c}\sim 0.6$, which is in agreement with the crossover threshold $\alpha_{\times}\sim 0.62$ predicted by the wave number selection and the Eckhaus instability. This is in direct contrast to the 2D disorder-to-order transition, which occurs at $\alpha\sim 0.28$ [3], and highlights the role of dimensionality on the phenomenology of the NRCH model.
@@ III.4 Resonances
A careful analysis of the $\rho_{D}$ vs. $\alpha$ plot reveals nontrivial features of the defect configurations. We find that $\rho_{D}$ shows sharp minima at certain values of $\alpha$ which we call “resonances” (see Fig. 4(a)). A visual inspection of the evolution of the defect configurations reveals that at resonances, the defects keep merging even at longer times resulting in smaller defect density. For further analysis, we compute the time it takes for the defect configurations to reach a steady state ($t_{\mathrm{st}}$), where $t_{\mathrm{st}}$ is defined as the time after which no defect merger event takes place and the defect density becomes constant. We find that for the majority of $\alpha$-space, defects attain a steady state relatively quickly, however in the resonance regions, $t_{\mathrm{st}}$ can be order of magnitude higher. Consistent with these observations, we also observe sharp rise in the nearest neighbour separation $d_{\mathrm{min}}$ in the resonance-space. Similar phenomenon, where $d_{\mathrm{min}}$ rises sharply and decreases gradually was also observed in the study of pairwise defect interactions in 2D [4]. It also highlights the rich phenomenology of defect interactions in the NRCH model.
IV Dirichlet and Neumann Boundary Conditions
We now consider the effects of varying boundary conditions on the phenomenology of the NRCH model. As discussed in Section II, in addition to the periodic boundary conditions, we consider Neumann and Dirichlet boundary conditions. Few comments are in order before we discuss the numerical results on N-BC and D-BC. Firstly, travelling waves of the form (3) are incompatible with both N-BC and D-BC, thus we do not expect states with complete polar order to emerge beyond $\alpha_{c}$. Secondly, for both N-BC and D-BC, polar order vanishes at the boundaries, i.e.,
@@ IV Dirichlet and Neumann Boundary Conditions
as $\phi$ or $\partial_{x}\phi$ are zero at the boundaries for D-BC and N-BC, respectively. Thus, the boundaries act as additional sources or sinks which make it possible to have any number of sources and sinks in the inner domain. In our analysis, we ignore the defects on the boundaries and only consider defects that are of the form described in Fig. 2 and are located inside the domain within $x\in(x_{\mathrm{pad}},L-x_{\mathrm{pad}})$. For both N-BC and D-BC, we heuristically choose $x_{\mathrm{pad}}=50$, which we find, filters the boundary defects well.
Figure 5: Comparison of $k_{l}$ for different $\alpha$ for $\alpha<\alpha_{c}$ for different boundary conditions. For a given $\alpha$, $k_{\infty}$ is higher for both N-BC and D-BC when compared to the $k_{\infty}$ for the P-BC. As a check, we also plot the $k_{\infty}$ obtained from Dedalus for the P-BC, which matches with our own simulations (dashed black line). $\alpha_{\times}=0.35$.
Figure 6: Phenomenology for the N-BC, which is similar to the P-BC for $\alpha<\alpha_{c}$. (a) Plot of average defect density $\rho_{D}$ versus $\alpha$, which matches very well with the statistics for P-BC. (b) Plot of average polar order. For $\alpha<\alpha_{c}$, we find $\bar{J}\sim 0$. For $\alpha>\alpha_{c}$, the system shows partial order. Near the transition point, we observe large intermittent fluctuations, which are reflected in the large fluctuations in $\bar{J}$ as well. (c) Comparison of different length scales. For $\alpha<\alpha_{c}$, $k_{\infty}$ is the dominant length scale and all other scales converge to it. Near the transition, we find large disagreements as well as large fluctuations and different length scales, which once again converge at larger $\alpha$.
Figure 7: Phenomenology for the D-BC. (a) Plot of average defect density $\rho_{D}$ versus $\alpha$, which is in agreement with the P-BC results for $\alpha>0.3$ For smaller $\alpha$, $\rho_{D}$ is significantly lower, and $\rho_{D}$ often vanishes for these $\alpha$. (b) Plot of average polar order. For $\alpha<\alpha_{c}$, when $\rho_{D}>0$, we once again find $\bar{J}\sim 0$. However, certain realizations for $\alpha\in(0.1,0.3)$ reach defect free configurations and thus show significant polar order (orange markers). For $\alpha>\alpha_{c}$, the phenomenology is similar to the N-BC; system shows partial order and near the transition point, we once again observe large fluctuations in $\bar{J}$. (c) Comparison of different length scales which show similar behaviour as N-BC.
Figure 8: Polar order $J$ for three different configurations with no defects, one defect, and multiple defects for $\alpha=0.2$ for D-BC. Markers show the location of defects inside the domain. Since, $|J|>0$ in the entire domain for defect-less configuration, $\bar{J}$ is large.
Figure 9: Kymographs of $J(x,t)$ for N-BC and D-BC for slightly above the transition point ($\alpha=0.70)$ and at higher values of nonreciprocity ($\alpha=1.0$). Closer to the transition point, we observe large transient and intermittent patches of polar order in the domain that keep emerging and vanishing. For large $\alpha$, both N-BC and D-BC show that the system develops two partitions with polar order in opposite directions. These partitions sustain over time but fluctuate around their mean value.
@@ IV.1 Defects below $\alpha_{c}$
Below $\alpha_{c}$, we once again find defect configurations. As shown in Fig. 5, similar to the P-BC, for both N-BC and D-BC, defects with a particular wave number are selected. However, wave number selection is affected by the presence of boundaries; for a given $\alpha$, the selected wave number is higher than that for the P-BC and a least-square fit yields $k_{\infty}=\alpha^{0.52}$. Thus, one expects that defect solutions will vanish around $\alpha\sim 1$. It is tempting to use the Eckhaus instability to arrive at a crossover threshold of nonreciprocity; $\alpha_{\times}\sim 0.33$ for N-BC and D-BC. However, we note that the instability analysis assumes travelling wave states, which are not permitted solutions for these boundary conditions. Our numerical analysis suggests that the transition occurs at $\alpha_{c}\sim 0.6$.
The phenomenology of the NRCH model with N-BC is qualitatively similar to the P-BC. For N-BC, the defect density $\rho_{D}$ agrees with the results from the P-BC (see Fig. 8(a)). Further, the average polar order vanishes for defect states. However, D-BC show a distinct phenomenology as compared to the two. For $\alpha\in(0.1,0.2)$, the system quickly attains a steady-state configuration, which can have any number of defects, including no defects at all. Thus the defect density shows large fluctuations for these values (see Fig. 8(a)). Since, the system can end up in defect-free or defect-laden states for D-BC (see Fig. 8), while computing the average polar order, we make a distinction between the two kind of states and find that as expected defect-laden configuration show negligible polar order. On the other hand, defect-free configurations exhibit large average polar order. For these states, one boundary acts as a source, and the other acts as a sink. Irrespective of the number of the defects remaining at the end, the system exhibits wave number selection and the selected wave number is the same. This is expected as the defect free state is achieved by consecutive merger of defects, and thus the wave number is selected prior to the formation of this state. For $\alpha\geq 0.3$, $\rho_{D}$ agrees with the results for P-BC.
@@ IV.2 Fluctuating polar order beyond $\alpha_{c}$
In the case of P-BC, a single wavenumber dominates the steady states, for defect configurations below $\alpha_{c}$ it is the selected wavenumber $k_{\infty}$, for the travelling states above $\alpha_{c}$, it is the wave number of the travelling wave. For N-BC and D-BC, we find that the intermittent patchy ordered states exhibit fluctuations at multiple wave numbers. To verify this, we compute various relevant scales using the structure factor $S(k,t)$: (i) The dominant mode $k_{\mathrm{d}}$ defined as the wavenumber at which $S(k)$ peaks, i.e, $k_{\mathrm{d}}=\mathrm{argmax}_{k}S(k)$, (ii) the coarsening mode $k_{\mathrm{c}}=\sum_{k}kS(k)/\sum_{k}S(k)$, and (iii) the integral mode $k_{\mathrm{i}}=\sum_{k}S(k)/\left(\sum_{k}S(k)/k\right)$. Here, $S(k,t)=|\psi(k)|^{2}$ is the structure factor, where $\psi(k,t)=\sum_{x}\psi(x,t)\exp(-ikx)$ is the Fourier transform of the windowed field $\psi(x,t)=w(x)\phi(x,t)$. Note that we apply the Hanning window $w(x)=\mathrm{sin}^{2}(\pi x/L)$ on the data for the N-BC and D-BC to avoid spurious amplitude errors arising from the aperiodicity. As shown in Fig. 8(c) and Fig. 8(c), $k_{\mathrm{d}}$, $k_{\mathrm{c}}$, and $k_{\mathrm{i}}$ are in excellent agreement with each other and match with $k_{\infty}$ for $\alpha<\alpha_{c}$, emphasizing that $k_{\infty}$ governs the dynamics for defect configurations. Slightly above $\alpha_{c}$, different modes disagree with each other, and exhibit large fluctuations, signifying that the transient states show fluctuations at multiple length scales. Consequently, it is not possible to choose a suitable normalization for the average polar order; i.e., a suitable choice for $q_{p}$ in Eq. 7 is not available for these states, which is also reflected in large fluctuations of $\bar{J}$ (see Fig. 8(b) and Fig. 8(b)) that were computed using $k_{\mathrm{c}}$.
For $\alpha>\alpha_{c}$, the system solves the incompatibility with the travelling waves by creating two domains with opposite polar order (see Fig. 9). The system boundary and the partition between the two domains with opposite polar order acts as sources or sinks of the polar order. The partition can form anywhere inside the domain and fluctuates slowly with time. In this case, the distribution function of $J(x,t)$ is primarily bimodal, where the strength of the two peaks depends the location of partition in the domain. As a consequence, we find large non zero polar order when the partition is closer to one of the domains, or small polar order if it close to the middle of the simulation domain. Finally, various scales, $k_{\mathrm{d}}$, $k_{\mathrm{c}}$, and $k_{\mathrm{i}}$, all converge to similar values.

### v2

@@ III.2 Wave number selection
As we show in Fig. 3, for a given $\alpha$, sources and sinks with a particular value of the asymptotic wave number $k_{\infty}$ are selected. $k_{\infty}$ increases monotonically with $\alpha$, and from a least-square fit, we obtain $k_{\infty}\sim\alpha^{0.6}$. Furthermore, $k_{\infty}$ is comparable to the wave number selected for targets in 2D. In Fig. 3(inset), we verify that $k_{\infty}$ for a single source-sink pair configuration is the same as the $k_{\infty}$ obtained from multi-defect configurations that evolved from disordered initial conditions. A monotonically increasing $k_{\infty}$ predicts a crossover value of non reciprocity $\alpha_{\times}$, such that when $\alpha\to\alpha_{\times}$, $k_{\infty}\to 1$, and $R_{\infty}\to 0$. This implies that the defect solutions will cease to exist for $\alpha\geq\alpha_{\times}$. Using a least-square fit, we find $\alpha_{\times}\sim 1.5$. However this prediction of $\alpha_{\times}$ is not entirely correct as it does not account for the Eckhaus instability. Since the plane waves generated by the defects are unstable for $k^{2}\geq 1/3$ [35, 8, 3, 41], we expect that the defect solutions will vanish for $k_{\infty}^{2}\geq 1/3$ which yields $\alpha_{\times}\sim 0.62$. Thus wavenumber selection and the Eckhaus instability restrict the defect solutions to the range $0<\alpha<0.62$. It is interesting to note that this threshold for the stability of the defects is directly connected to the onset of the emergence of global polar order (see below).
@@ III.3 Defect density and the Global polar order
We now focus on the statistical properties of the multi-defect configurations. Using the fact that $J=0$ at a source or a sink, we define the 1D defect density $\rho_{D}$ simply as the number of zeros of $J$ per unit length. In Fig. 4(a), we plot $\rho_{D}$ in the steady state for different values of $\alpha$. Since $\alpha=0$ is the equilibrium limit of the model, $\rho_{D}$ vanishes as $\alpha\to 0$. Furthermore, $\rho_{D}$ vanishes above $\alpha_{c}\sim 0.6$ and instead we find travelling waves. This marks the onset of disorder-to-order transition that was also observed in 2D [3]. To better analyse the transition, in Fig. 4(b), we plot the average polar order $\bar{J}$ versus $\alpha$, in the steady state. $\bar{J}$ is defined as
@@
where $\left$$\ldots\right$> implies averaging over space and time in the steady-state, and we use the dominant wave number of the patterns $q_{p}$ for normalization. For travelling waves, $J=R^{2}k=k-k^{3}$, and $q_{p}=k$, which sets $\bar{J}=1$ for all travelling waves, irrespective of their wave number. For defects, $q_{p}=k_{\infty}$ is a suitable choice. Since defects are sources or sinks of travelling waves in both the directions, we expect $\bar{J}$ to vanish in the presence of defects. For $\alpha<\alpha_{c}$, $\bar{J}$ is close to zero and for $\alpha>\alpha_{c}$, consistent with the travelling states, $\bar{J}\sim 1$.
For very small $\alpha$, or for $\alpha\lesssim\alpha_{c}$, $\rho_{D}$ is small, which implies that there are only a handful of defects in the domain. This can lead to finite size effects and large fluctuations that are reflected in the plot of $\bar{J}$. However, these effects become smaller as the domain size increases.
Another consequence of the finite domains is that in the vicinity of $\alpha_{c}$ (grey shaded region), the steady state varies across the independent realizations of the initial conditions. Some of the simulations transition to travelling waves while others show defects. Thus, to find a better estimate of $\alpha_{c}$, we compute the fraction of simulations $(F_{D})$ that show defect states (see Fig. 4(b,inset)). For $\alpha\lesssim 0.4$, $F_{D}=1$ which implies that the entire ensemble shows defect states. For $\alpha>0.6$, all the simulations for the largest domains end up in travelling wave states. From the plots of $\rho_{D}$ and $\bar{J}$, we conclude that the disorder-to-order transition in 1D occurs at $\alpha_{c}\sim 0.6$, which is in agreement with the crossover threshold $\alpha_{\times}\sim 0.62$ predicted by the wave number selection and the Eckhaus instability.
To understand this result, we note that the travelling wave solutions of NRCH are linearly stable at all values of $\alpha$ [8], and therefore, the only reason they might not be adopted at relatively smaller values of $\alpha$ is the emergence of relatively stable defects that form an arrested defect network [3, 4]. Once these defects are destabilized (beyond $\alpha_{\times}\sim 0.62$) due to Eckhaus instability, the only stable solutions that remain will be the travelling waves. While the above argument, which supports $\alpha_{c}\simeq\alpha_{\times}$, provides a simple and comprehensive explanation for the 1D phase diagram, we note that this scenario does not hold for the disorder-to-order transition in 2D, where $\alpha_{c}\sim 0.28$ is much smaller than $\alpha_{\times}\sim 0.58$ [3]. This suggests that in 2D there must be other effects (not present in 1D) that can destabilize the defects before the Eckhaus instability occurs. These effects could originate from the role of geometry of the domains, the existence of different types of defects, etc, and highlight the role of dimensionality in the phenomenology of the NRCH model.
@@ III.4 Resonances
A careful analysis of the $\rho_{D}$ versus $\alpha$ plot reveals nontrivial features of the defect configurations. We find that $\rho_{D}$ shows sharp minima at certain values of $\alpha$ which we call “resonances” (see Fig. 4(a)). A visual inspection of the evolution of the defect configurations reveals that at resonances, the defects keep merging even at longer times resulting in smaller defect density. For further analysis, we compute the time it takes for the defect configurations to reach a steady state ($t_{\mathrm{st}}$), where $t_{\mathrm{st}}$ is defined as the time after which no defect merger event takes place and the defect density becomes constant. We find that for the majority of $\alpha$-space, defects attain a steady state relatively quickly, however in the resonance regions, $t_{\mathrm{st}}$ can be order of magnitude higher (see Fig. 5(a)). Consistent with these observations, we also observe sharp rise in the nearest neighbor separation $d_{\mathrm{min}}$ for the resonances (see Fig. 5(b)). Similar phenomenon, where $d_{\mathrm{min}}$ rises sharply and decreases gradually was also observed in the study of pairwise defect interactions in 2D [4].
IV Effect of Boundary Conditions
We now consider the effects of varying boundary conditions on the phenomenology of the NRCH model. As discussed in Section II, in addition to the periodic boundary conditions, we consider Neumann and Dirichlet boundary conditions. Traveling waves of the form (6) are incompatible with both N-BC and D-BC, thus we do not expect states with complete polar order to emerge beyond $\alpha_{c}$. Further, for both N-BC and D-BC, polar order vanishes at the boundaries, i.e.,
@@ IV Effect of Boundary Conditions
as either $\phi$ or $\partial_{x}\phi$ vanish at the boundaries. Thus, the boundaries act as additional sources or sinks and in our analysis, we ignore the defects on the boundaries and only consider defects that are of the form described in Fig. 2 and are located inside the domain within $x\in(x_{\mathrm{pad}},L-x_{\mathrm{pad}})$. For both N-BC and D-BC, we heuristically choose $x_{\mathrm{pad}}=50$, which we find, filters the boundary defects very well.
Figure 6: Comparison of $k_{\infty}$ for different $\alpha$ for $\alpha<\alpha_{c}$ for different boundary conditions. For a given $\alpha$, $k_{\infty}$ is higher for both N-BC and D-BC when compared to the $k_{\infty}$ for the P-BC. As a check, we also plot the $k_{\infty}$ obtained from Dedalus for the P-BC, which matches with our own simulations (dashed black line).
Figure 7: Phenomenology for the N-BC (a-c) and D-BC (d-f). For $\alpha<\alpha_{c}$, the phenomenology is similar to the P-BC. (a, d) Plot of average defect density $\rho_{D}$ versus $\alpha$, which is consistent with the statistics for P-BC within the error bars. (b, e) Plot of the average polar order. For $\alpha<\alpha_{c}$, we find $\bar{J}\sim 0$. For $\alpha>\alpha_{c}$, the system shows partial order. Near the transition point, we observe large intermittent fluctuations, which are reflected in the large fluctuations in $\bar{J}$ as well. (c, f) Comparison of different length scales. For $\alpha<\alpha_{c}$, $k_{\infty}$ is the dominant length scale and all other scales converge to it. Near the transition, we find large disagreements as well as large fluctuations and different length scales, which once again converge at larger $\alpha\sim 0.9$.
@@ IV.1 Defects below $\alpha_{c}$
Below $\alpha_{c}$, we once again find defect configurations. As shown in Fig. 6, similar to the P-BC, for both N-BC and D-BC, defects with a particular wave number are selected. However, wave number selection is affected by the presence of boundaries; for a given $\alpha$, the selected wave number is higher than that for the P-BC and a least-square fit yields $k_{\infty}=\alpha^{0.52}$. Thus, one expects that defect solutions will vanish around $\alpha\sim 1$. It is tempting to use the Eckhaus instability to arrive at a crossover threshold of nonreciprocity; $\alpha_{\times}\sim 0.33$ for N-BC and D-BC. However, we note that the instability analysis assumes travelling wave states, which are not permitted solutions for these boundary conditions. Our numerical analysis suggests that the defect states persist up to $\alpha_{c}\sim 0.6$.
The phenomenology below $\alpha_{c}$ is qualitatively similar to the P-BC. The defect density $\rho_{D}$ agrees with the results from the P-BC (see Fig. 7(a, d)). Further, the average polar order vanishes for defect states.
@@ IV.2 Fluctuating polar order beyond $\alpha_{c}$
In the case of P-BC, a single wavenumber dominates the steady states, for defect configurations below $\alpha_{c}$ it is the selected wavenumber $k_{\infty}$, for the travelling states above $\alpha_{c}$, it is the wave number of the travelling wave. For N-BC and D-BC, we find that the intermittent patchy ordered states exhibit fluctuations at multiple wave numbers. To verify this, we compute various relevant scales using the structure factor $S(k,t)$:
  1. (i)
    The dominant mode $k_{\mathrm{d}}$ defined as the wavenumber at which $S(k)$ peaks, i.e, $k_{\mathrm{d}}=\mathrm{argmax}_{k}S(k)$.
  2. (ii)
    The coarsening mode $k_{\mathrm{c}}=\sum_{k}kS(k)/\sum_{k}S(k)$
  3. (iii)
    The integral mode $k_{\mathrm{i}}=\sum_{k}S(k)/\left(\sum_{k}S(k)/k\right)$.
Here, $S(k,t)=|\psi(k)|^{2}$ is the structure factor, where $\psi(k,t)=\sum_{x}\psi(x,t)\exp(-ikx)$ is the Fourier transform of the windowed field $\psi(x,t)=w(x)\phi(x,t)$. Note that we apply the Hanning window $w(x)=\mathrm{sin}^{2}(\pi x/L)$ on the data for the N-BC and D-BC to avoid spurious amplitude errors arising from the aperiodicity.
As shown in Fig. 7(c,f), $k_{\mathrm{d}}$, $k_{\mathrm{c}}$, and $k_{\mathrm{i}}$ are in excellent agreement with each other and match with $k_{\infty}$ for $\alpha<\alpha_{c}$, emphasizing that $k_{\infty}$ governs the dynamics for defect configurations. Slightly above $\alpha_{c}$, different modes disagree with each other, and exhibit large fluctuations, signifying that the transient states show fluctuations at multiple length scales. Consequently, it is not possible to choose a suitable normalization for the average polar order; i.e., a suitable choice for $q_{p}$ in Eq. 10 is not available for these states, which is also reflected in large fluctuations of $\bar{J}$ (see Fig. 7(b,d)) where we have used $q_{p}=k_{\mathrm{c}}$.
For $\alpha>\alpha_{c}$, the system solves the incompatibility with the travelling waves by creating two domains with waves travelling in the opposite direction (see Fig. 8). These domains are separated by a partition where $|\phi|$ vanishes. and in addition to the domain boundaries, it acts as a source or sink for the travelling waves. The partition can form anywhere inside the domain and fluctuates slowly with time. The distribution function of $J(x,t)$ is primarily bimodal, where the strength of the two peaks depends on the location of partition in the domain. As a consequence, we find large non zero polar order when the partition is closer to one of the domains, or small polar order if it is close to the middle of the simulation domain. Finally, various scales, $k_{\mathrm{d}}$, $k_{\mathrm{c}}$, and $k_{\mathrm{i}}$, all converge to similar values.
Figure 8: Kymographs of $J(x,t)$ for N-BC and D-BC for $\alpha=0.70$ (slightly above $\alpha_{c}$) and for $\alpha=1.0$. Closer to the transition point, we observe large transient and intermittent patches of polar order in the domain that keep emerging and vanishing. For large $\alpha$, both N-BC and D-BC show that the system develops two partitions with polar order in opposite directions. The position of the partition fluctuates with time.